OpenStudy (anonymous):

What is the value of the expression 7^–3? 1/343 ? 3 exponent

jimthompson5910 (jim_thompson5910):

very good, \[\Large 7^{-3} = \frac{1}{7^3}\] \[\Large 7^{-3} = \frac{1}{343}\]

OpenStudy (xxdragoonxx):